# PointNSP: Autoregressive 3D Point Cloud Generation with Next-Scale Level-of-Detail Prediction

### v2

@@ 3.2 Multi-Scale LoD Representation
Here, $\mathcal{L}_{\text{CD}}$ denotes the Chamfer Distance loss, and $\mathcal{L}_{\text{EMD}}$ denotes the Earth-Mover Distance loss. These two terms are commonly used in prior works to measure the distance between point clouds from complementary perspectives. The stop-gradient operation, $sg[\cdot]$, ensures that the latent features $f$ used for reconstruction remain consistent with the quantized latent vectors $\mathbf{z}$.
Figure 4: Illustration of training a multi-scale 3D point cloud causal transformer with intermediate decoding, $\operatorname{upsampling}$, position-aware soft masks, and block-wise causal masks.
3.3 Autoregressive Transformer for Next-scale LoD Prediction
The next step is to train an autoregressive transformer on the multi-scale token sequence $Q=(q_{1},\dots,q_{K})$ obtained from the previous stage. Due to the strong local-geometry inductive bias inherent in 3D structures and the relatively smaller dataset size compared to 2D images, a naive causal transformer struggles to efficiently capture both intra-scale and inter-scale interactions. Consequently, it becomes necessary to embed geometric information into the transformer design, though this introduces additional challenges due to the unordered nature of point clouds.
Inter-Scale Interaction Modeling.
Inter-scale interactions across all scales $(q_{1},\dots,q_{K})$ are essential for the model to capture relationships between scales and to generate the next LoD conditioned on the preceding ones. The model is guided by a key principle: tokens at scale $\hat{q}_{k}$ are restricted to attend only to tokens from the preceding scales $\hat{q}_{1},\dots,\hat{q}_{k-1}$. Within each scale, however, all tokens $q_{k}=(q^{1}_{k},\dots,q^{s_{k}}_{k})$ are allowed to fully attend to one another, ensuring that the model interprets them as complete shapes at the corresponding resolution. To enforce this constraint, we construct the causal mask $\mathbf{M}$ as a block-diagonal matrix, where each diagonal block $\mathbf{M}_{k}$ of size $s_{k}\times s_{k}$ is fully unmasked: $\mathbf{M}=\operatorname{diag}[\mathbf{M}_{1},\mathbf{M}_{2},...,\mathbf{M}_{K}]$. This design allows bidirectional dependencies within each scale while preserving a lower-triangular structure across scales, thereby preventing information leakage from future tokens. To further distinguish scales, each token is assigned a scale embedding $\mathbf{s}^{i}_{k}\in\mathbb{R}^{d}$, implemented as a one-hot vector over the $K$ scales and shared among tokens within the same scale. Some algorithmic details are explained in the Appendix B.
Intra-Scale Interaction Modeling.
Capturing intra-scale interactions within the token set $q_{k}=(q^{1}_{k},\dots,q^{s_{k}}_{k})$ is crucial for the model to effectively understand shapes across different resolutions. An important geometric prior is the relative distance between pairs of point tokens. Incorporating this information into intra-scale bidirectional attention allows tokens to place greater weight on nearby neighbors rather than starting from uniform attention. Specifically, we compute the position-aware soft masking matrix $\mathbf{M}^{p}_{k}\in\mathbb{R}^{s_{k}\times s_{k}}$, which is derived from the coordinate-based positional embedding matrix $\mathbf{P}_{k}\in\mathbb{R}^{s_{k}\times d}$, as follows:
$\mathbf{M}^{p}_{k}$ is a symmetric matrix, where each entry $\mathbf{M}^{p}_{k}[i,j]\in(0,1)$ encodes the relative position between points $i$ and $j$. This raises a key question: how can we derive the positional embedding $\mathbf{P}_{k}$ for each scale when no explicit 3D geometry is available at this stage? Our solution is an intermediate-structure decoding strategy. Specifically, for each scale $k$, we apply the decoder $D(\cdot)$ to reconstruct the intermediate structure $\mathbf{\hat{X}_{k-1}}$ using all tokens predicted up to step k1k-1:
$\mathbf{\hat{X}_{k-1}}=D(\sum_{m=1}^{k-1}\phi_{m}(\operatorname{upsampling}(\mathbf{\hat{z}}_{m},s_{m}))),\ \{\mathbf{\hat{z}}_{1},\dots,\mathbf{\hat{z}}_{k-1}\}=\operatorname{lookup}(Z,\{\hat{q}_{1},\dots,\hat{q}_{k-1}\}).$ (9)
$\mathbf{\hat{X}_{k-1}}\in\mathbb{R}^{s_{k-1}\times 3}$ serves as the input coordinate information for the next scale $k$ prediction. The positional embedding is derived as an absolute positional encoding based on the 3D coordinates $\mathbf{\hat{X}_{k-1}}$ using trigonometric functions (e.g. $\sin$ and $\cos$). Constructing the encoding directly from 3D coordinates, rather than point indices, preserves permutation equivariance. The detailed derivation of $\mathbf{P}$ is provided in the Appendix B. Note that any positional encoding based on token indices should not be applied, as it would violate the permutation-equivariant property.
The prediction of each token $\hat{q}_{k}^{i}$ is evaluated using the cross-entropy (CE) loss $\mathcal{L}^{i}_{k}$. We first compute intra-scale loss $\mathcal{L}_{k}=\frac{1}{s_{k}}\sum_{i=1}^{s_{k}}\mathcal{L}^{i}_{k}$ and then compute inter-scale loss $\mathcal{L}_{\text{total}}=\frac{1}{K}\sum_{k=1}^{K}\mathcal{L}_{k}$. The second stage training architecture is illustrated in Figure 4. We provide a theoretical analysis of the permutation invariance of PointNSP ’s distribution modeling in Appendix A.

### v3

@@ 3.3 Autoregressive Transformer for Next-Scale LoD Prediction
LION Diffusion 67.41 61.23 53.70 52.34 53.41 51.14 58.17 54.90
TIGER Diffusion 67.21 56.26 54.32 51.71 54.12 50.24 58.55 52.74
[…]
Table 1: Performance under the standard $2048$-point setup on ShapeNet is reported for two dataset splits: the top corresponds to the conventional random split, and the bottom corresponds to the LION split [LION]. The best results are highlighted in bold with a green bar, and the second-best results are underlined.
Figure 4: Visualization of generation results compared with baseline models. PointNSP produces high-quality and diverse 3D point clouds.
@@ 4 Experiments
  Model Dense Point Cloud Generation (8192 pts) Many-Class Generation (55-Class)
Airplane Chair Car Airplane Chair Car
CD $\downarrow$ EMD $\downarrow$ CD $\downarrow$ EMD $\downarrow$ CD $\downarrow$ EMD $\downarrow$ CD $\downarrow$ EMD $\downarrow$ CD $\downarrow$ EMD $\downarrow$ CD $\downarrow$ EMD $\downarrow$
  PVD 69.77 69.98 52.56 51.33 54.19 46.55 97.53 99.88 88.37 96.37 89.77 94.89
TIGER 68.48 60.24 51.87 51.85 53.78 44.12 83.54 81.55 57.34 61.45 65.79 57.24
PointGPT 69.29 61.56 52.43 52.09 54.97 45.20 94.94 91.73 71.83 79.00 89.35 87.22
LION 68.95 60.38 53.76 53.45 54.98 44.67 86.30 77.04 66.50 63.85 64.52 54.21
PointNSP-s (ours) 67.84 58.25 51.26 51.81 52.40 43.76 78.95 68.84 58.79 55.10 59.97 52.89
[…]
Table 2: Comparison of dense point cloud generation (left) and many-class generation (right). CD and EMD metrics ($\downarrow$) are reported.
4.1 Experimental Setup
Datasets & Metrics.
In line with prior studies, we adopt ShapeNetv2, pre-processed by PointFlow [PointFlow], as our primary dataset. Each shape is globally normalized and uniformly sampled to $2048$ points (standard setting). Experiments under this setting are conducted on two data splits: the standard random split and the LION split [LION, Tiger]. We further evaluate a dense generation setting with $8192$ points (dense setting). Following established benchmarks [PVD, LION], we use the 1-nearest neighbor (1-NN) accuracy [1-NN] as our primary evaluation metric, which effectively captures both the quality and diversity of generated point clouds—a score near $50\%$ indicates strong performance [PointFlow]. The 1-NN distance matrix is computed using two widely adopted distance measures: Chamfer Distance (CD) and Earth Mover’s Distance (EMD). We also report mean CD and mean EMD, averaged across all object categories. To assess efficiency, we record the total training time in GPU hours and report both inference time and parameter count.
[…]
$downarrow Training\downarrow Sampling\downarrow$
6 PointGPT 185 5.32 296 10.56 46
5 PVD 142 29.9 201 58.1 45
4 LION 550 31.2 610 59.5 60
3 TIGER 164 23.6 320 42.1 55
[…]
1 PointNSP-m (ours) 178 3.59 190 5.48 32
Table 3: Training time (in GPU hours, averaged over three categories), sampling time (in seconds, averaged over samples), and model size (in millions of parameters). Ranked by generation quality on $2048$ and $8192$ settings.
4.2 Single-Class Generation: Standard & Dense
Results.
To comprehensively assess the performance of PointNSP, we compare it with several strong baseline models. Specifically, we include SoTA diffusion-based approaches—ShapeGF [ShapeGF], DPM [DPM], PVD [PVD], LION [LION], and TIGER [Tiger]—as well as leading autoregressive models, including PointGrow [PointGrow], CanonicalVAE [pointvqvae], and PointGPT [PointGPT]. We report results for two variants of our model, PointNSP-s and PointNSP-m, corresponding to small and medium parameter configurations, respectively. Note that we only compare against models with publicly available implementations. Quality. As shown in Table 1, PointNSP consistently surpasses all baseline methods on both the conventional and LION data splits under the standard $2048$-point setting, establishing new SoTA generation quality. Notably, even the lightweight PointNSP-s achieves highly competitive results. Under the dense $8192$-point setting, Table 2 (Left) further demonstrates that PointNSP outperforms strong baselines by an increasingly larger margin. Efficiency. Table 3 reports the training time, sampling time, and model size for several strong baseline models from Table 1, including PVD, LION, TIGER, and PointGPT. Among these, PointNSP-s achieves the shortest training time, fastest sampling speed, and highest parameter efficiency, while still delivering competitive performance. PointNSP-m attains state-of-the-art generation quality with the second-best efficiency, slightly trailing its lighter counterpart. Under the dense $8192$-point setting, PointNSP-m requires only about half as many parameters as leading diffusion-based models while maintaining superior performance. Collectively, these results highlight the strong scalability potential of PointNSP.
4.3 Many-Class Unconditional Generation
Beyond the Single-Class generation setting, we further evaluate our model on the more challenging Many-Class generation task introduced by LION [LION], aiming to assess its ability to generalize across diverse object categories. Specifically, we train PointNSP and selected strong baseline models without using class conditioning over $55$ distinct categories from ShapeNet. This poses a challenge for the model to capture multi-modal and structurally complex geometric patterns across categories. As shown in Table 2 (Right), PointNSP significantly outperforms all strong baselines. We provide visualizations of diverse shape classes in Appendix 10.
4.4 Point Cloud Completion & Upsampling
We evaluate PointNSP on two key downstream tasks: point cloud completion and upsampling. For completion, we follow the experimental setup of PVD [PVD]. For upsampling, we use a factor of $r=2$, increasing the input point clouds from $1024$ to $2048$ points. As shown in Table 5, PointNSP consistently achieves the best performance on point cloud completion. For the upsampling task, it outperforms all selected baselines across both metrics, highlighting its effectiveness on diverse downstream applications and its potential as a foundation model. Visualization results for these two tasks are presented in Figure 6 and Appendix 10.
4.5 Ablation and Analysis
Position Mask Upsampling FPS Stochasticity TE Mean CD $\downarrow$ Mean EMD $\downarrow$
Voxel PU-Net
SE 64.25 60.53
SE 63.86 59.95
SE+A-PE 62.19 58.23
SE+A-PE 63.05 58.47
L-PE 63.22 59.71
A-PE 62.12 60.02
A-PE 61.28 57.32
SE+L-PE 60.62 57.34
[…]
Table 4: Training time (in GPU hours, averaged over three categories), sampling time (in seconds, averaged over samples), and model size (in millions of parameters). Ranked by generation quality on $2048$ and $8192$ settings.
  Category Point Cloud Completion Point Cloud Upsampling
Model CD $\downarrow$ EMD $\downarrow$ Model CD $\downarrow$ EMD $\downarrow$
  Airplane SoftFlow 40.42 11.98 PVD 73.56 71.65
PointFlow 40.30 11.80 TIGER 71.65 59.94
DPF-Net 52.79 11.05 PointGPT 72.11 60.12
PVD 44.15 10.30 LION 70.41 59.65
[…]
Chair SoftFlow 27.86 32.95 PVD 53.81 64.61
PointFlow 27.07 36.49 TIGER 52.80 52.98
DPF-Net 27.63 33.20 PointGPT 53.75 53.21
PVD 32.11 29.39 LION 53.98 54.33
[…]
Car SoftFlow 18.50 27.89 PVD 58.95 48.43
PointFlow 18.03 28.51 PointGPT 57.26 47.85
DPF-Net 13.96 23.18 TIGER 57.90 48.01
PVD 17.74 21.46 LION 57.14 47.56
[…]
Table 5: Comparison on partial shape completion (left) and point cloud upsampling task (right).
We conduct comprehensive ablation studies to assess the impact of various architectural components and training strategies. First, we compare two upsampling strategies: voxel-based representations and PU-Net. Our experiments show that PU-Net consistently outperforms voxel-based upsampling, owing to its permutation-equivariant design. Second, we evaluate the effectiveness of the position-aware masking strategy, which significantly boosts model performance. Third, we analyze the contribution of each embedding in the token embedding layer, with results highlighting the notable impact of the scale embedding. We also compare learnable positional encoding (L-PE) with absolute PE (A-PE), finding that A-PE yields better performance. Finally, we examine the impact of $\operatorname{FPS}$ stochasticity, which stems from the inherent randomness of the $\operatorname{FPS}$ algorithm. Compared to a variant with a fixed initialization that yields deterministic downsampling, our results show that incorporating FPS stochasticity consistently improves model generalization. These findings are summarized in Table 4, with additional ablations—such as varying the number of scales—presented in Appendix 8.
5 Conclusions
In this work, we introduce PointNSP, a novel autoregressive framework for high-quality 3D point cloud generation. Unlike previous approaches, PointNSP employs a coarse-to-fine strategy that captures the multi-scale level of detail (LoD) of 3D shapes while preserving global structural coherence, rather than decomposing generation into local predictions. Our method consistently surpasses existing point cloud generative models in quality, while achieving substantial gains in parameter efficiency, training efficiency, and inference speed. Looking ahead, we aim to extend PointNSP toward foundation-scale models trained on large-scale 3D datasets as a promising future direction.
[…]
Supplementary Material
6 Permutation Invariance of Probability Distribution
Here we show why the distribution $p(\mathbf{X}_{1},\mathbf{X}_{2},\dots,\mathbf{X}_{K})=\prod_{k=1}^{K}p(\mathbf{X}_{k}|\mathbf{X}_{k-1},\dots,\mathbf{X}_{2},\mathbf{X}_{1})$ in Eq. 3 preserves the permutation invariance property $p(\pi(\mathbf{x}_{1},\dots,\mathbf{x}_{N}))=p(\mathbf{x}_{1},\dots,\mathbf{x}_{N}),\forall\ \pi\in S_{N}$ in Eq. 2.
By definition, the joint distribution factorizes as:
$p(\mathbf{X}_{1},\mathbf{X}_{2},\dots,\mathbf{X}_{K})=p(\mathbf{X}_{1})p(\mathbf{X}_{2}|\mathbf{X}_{1})$ (10)
$\dots p(\mathbf{X}_{K}|\mathbf{X}_{1},\dots\mathbf{X}_{K-1}).$
$\pi=(\pi_{1},\pi_{2},\dots,\pi_{K}),\hskip 28.80008pt\pi_{k}\in S_{s_{k}}.$ (11)
Then we aim to prove that:
$p(\pi_{1}(\mathbf{X}_{1}),\dots,\pi_{K}(\mathbf{X}_{K}))=p(\mathbf{X}_{1},\dots\mathbf{X}_{K}).$ (12)
Recall that $\operatorname{FPS}$ is permutation-invariant, the resulting LoD sampling sequence $(\mathbf{X}_{1},\dots,\mathbf{X}_{K})$ is permutation-invariant (its inherent stochasticity operates at the set level and is independent of point ordering). The core components of PointNSP —namely the feature encoder $\mathcal{E}(\cdot)$, $\operatorname{upsampling}$, $\operatorname{query}$, and the decoder $D(\cdot)$—are all permutation-equivariant, ensuring that each output feature remains aligned with its corresponding input point. Therefore, the mapping between $\mathbf{X}_{k}$ and conditioning context shapes $(\mathbf{X}_{1},\dots,\mathbf{X}_{k-1})$ remains unchanged with respect to any global permutation $\pi$. Then, for any permutation $\pi_{k}$ of points within $\mathbf{X}_{k}$:
$p(\pi_{k}(\mathbf{X}_{k})|\pi_{1}(\mathbf{X}_{1}),\dots\pi_{k-1}(\mathbf{X}_{k-1}))$ (13)
$=p(\mathbf{X}_{k}|\mathbf{X}_{1},\dots\mathbf{X}_{k-1}).$
This holds for each $k=1,\dots,K$. Then the joint distribution under these permutations is
$p(\pi_{1}(\mathbf{X}_{1}),\dots,\pi_{K}(\mathbf{X}_{K}))$ (14)
$=\prod_{k=1}^{K}p(\pi_{k}(\mathbf{X}_{k})|\pi_{1}(\mathbf{X}_{1}),\dots\pi_{k-1}(\mathbf{X}_{k-1}))$
$=\prod_{k=1}^{K}p(\mathbf{X}_{k}|\mathbf{X}_{k-1},\dots,\mathbf{X}_{2},\mathbf{X}_{1})$
$=p(\mathbf{X}_{1},\dots,\mathbf{X}_{K})$
Figure S1: Illustration of different LoD sequences for a single shape. Owing to the inherent stochasticity of the sampling strategy in Eq. 17, our method naturally produces diverse causal LoD sequences for the same 3D shape across training epochs. At each scale, these variations encourage broader spatial coverage of the underlying geometry, thereby enhancing learning robustness.
Consequently, the autoregressive factorization preserves permutation invariance: permuting the points within any scale does not change the resulting joint probability:
7 Algorithm Details
LoD Sequence Sampling.
This step is a core component of PointNSP training, and we provide additional details here to eliminate any potential ambiguity. If one were to directly apply the original image-based VAR framework [VAR] to 3D point clouds, latent features for sampled LoD sequences would be obtained via
$\mathbf{f}_{k}=\operatorname{FPS}(\mathbf{f}^{k-2}-\tilde{\mathbf{f}}_{k-1}),\mathbf{f}_{1}=\operatorname{FPS}(\mathbf{f}^{0}).$ (16)
However, this strategy is unsuitable for 3D point clouds. Applying $\operatorname{FPS}$ directly in latent space implicitly assumes that latent distances form a meaningful metric, yet these distances need not correlate with underlying geometric structure. As a result, sampling in latent space cannot reliably preserve geometric uniformity or spatial coverage. To address this, rather than iteratively applying $\operatorname{FPS}$ in latent space, we apply $\operatorname{FPS}$ in Euclidean 3D space in a fine-to-coarse manner:
$\mathbf{X}_{k-1}=\operatorname{FPS}(\mathbf{X}_{k}),\mathbf{X}_{K}=\mathbf{X}.$ (17)
This produces a geometrically consistent LoD sequence $(\mathbf{X}_{1},\dots,\mathbf{X}_{K})$ from coarse to fine, with guaranteed point correspondences $\mathbf{X}_{1}\subsetneq\dots\subsetneq\mathbf{X}_{K-1}\subsetneq\mathbf{X}_{K}$. We then obtain latent features by querying the indices mapping each $\mathbf{X}_{k}$ back to the original point set:
$\mathbf{f}_{k}=\operatorname{query}(\mathbf{f}^{k-2}-\tilde{\mathbf{f}}_{k-1},\mathbf{X}_{k}),\ \mathbf{f}_{1}=\operatorname{query}(\mathbf{f}^{0},\mathbf{X}_{1}),$ (18)
where $\mathbf{f}^{0}=\mathcal{E}(\mathbf{X})$. This is Eq. 4. Here we need to emphasize that the LoD sequence $(\mathbf{X}_{1},\dots,\mathbf{X}_{K})$ is constructed in advance following Eq. 17 and the Line 5 ($\mathbf{X}_{k}=\operatorname{FPS}(\mathbf{X},s_{k})$) in Algorithm 1 denotes that $\mathbf{X}_{k}$ with scale $s_{k}$ is obtained starting from $\mathbf{X}$. Note that it does not mean the $\mathbf{X}_{k}$ is obtained by directly applying $\operatorname{FPS}$ on the original $\mathbf{X}$ as this would break the point correspondence across scales (i.e. $\mathbf{X}_{i}\cap\mathbf{X}_{j}\not=\mathbf{X}_{i},i<j$). Since $\operatorname{FPS}$ is inherently stochastic due to its random initialization, we can naturally obtain diverse LoD sequences for each shape $\mathbf{X}$ across training epochs. This variability is desirable, as the refinement path need not be unique; sampling multiple paths improves spatial coverage, as discussed in Section 3.2. For example, for epoch 0 and epoch $1$, we may obtain two LoD sequences for shape $\mathbf{X}$:
$(\mathbf{X}^{0}_{1},\dots,\mathbf{X}^{0}_{K}),\ (\mathbf{X}^{1}_{1},\dots,\mathbf{X}^{1}_{K}).$ (19)
At each scale, $\mathbf{X}^{0}_{k}$ and $\mathbf{X}^{1}_{k}$ represent the same underlying surface but with different point coverages, which helps improve learning robustness and generalization. We illustrate this LoD sequence sampling issue in Figure S1.
Coordinate-based Positional Encoding.
[…]
$p$ $=\lambda^{2}*z_{i}+\lambda*y_{i}+x_{i}$ (20)
$\mathbf{P}_{k}(p,2i)=\sin(\frac{p}{10000^{\frac{2i}{D}}})$ (21)
$\mathbf{P}_{k}(p,2i+1)=\cos(\frac{p}{10000^{\frac{2i}{D}}}),$ (22)
[…]
Intra-Scale Token Embedding.
Instead of simply adding all token embeddings together, the real implementation is analogous to Llama [Llama3] by adding positional embedding and scale embedding to query and key vectors. Specifically, we retrieve the codebook embedding $\mathbf{z}_{k}$ for each scale token $q_{k}$ and then upsampled from input scale to output scale $(s_{k}\rightarrow s_{k+1})$: $\mathbf{z}_{k}=\operatorname{upsampling}(\mathbf{z}_{k})\in\mathbb{R}^{s_{k+1}\times d}$ following the Eq. 6. Positional embedding $\mathbf{p}^{i}_{k}=\mathbf{P}_{k}[i]$ for each token $q_{k}^{i}$ is derived with from the decoded intermediate structure $\mathbf{X}_{k}$ as described in Eq. 9. During inference stage, the intermediate structure $\hat{\mathbf{X}}_{k}$ is decoded using the predicted token $\hat{q}_{k}$ instead. Additionally, the model needs to know which scale that each token belongs to. Therefore, $\mathbf{s}_{k}$ is a simple one-hot embedding $\mathbf{s}_{k}=\operatorname{one-hot-embedding}(k)$ out of total $K$ scales. Tokens from the same scale $k$ share the same scale embedding $\mathbf{s}_{k}$ ($\mathbf{s}^{i}_{k}=\mathbf{s}^{j}_{k}$ for $q_{k}^{i}$, $q_{k}^{j}$). For all input scale tokens, we add both the positional embedding $\mathbf{p}^{i}_{k}$ and the scale embedding $\mathbf{s}^{i}_{k}$ to the query $\mathbf{u}_{k}^{i}$ and key vectors $\mathbf{v}_{k}^{i}$ derived in the attention mechanism:
where $\mathbf{W}_{\mathbf{U}}$ and $\mathbf{W}_{\mathbf{V}}$ are projection matrices for queries and keys respectively. Together with value vectors, these vectors will be fed to the block-wise causal transformer for next-scale token prediction.
8 Hyperparameters & Reproducibility Settings
Hyperparameters & Reproducibility Settings.
Specifically, we set the learning rate $3e^{-4}$ and the batch-size 32. We perform all the experiments on a workstation with Intel Xeon Gold 6154 CPU (3.00GHz) and 8 NVIDIA Tesla V100 (32GB) GPUs. We use an AdamW optimizer with an initial learning rate of $10^{-4}$ for VQVAE training and $10^{-3}$ for autoregressive transformer training respectively. For upsampling and completion experiments, we follow the experimental settings of PVD [PVD]. For many-class generation, we mainly inherit the experimental setting from LION [LION].
Hyperparameter Value
# PVCNN layers 4
# PVCNN hidden dimension 1024
# PVCNN voxel grid size 32
# MLP layers 6
# Attention dimension 1024
# Attention head 32
Optimizer AdamW
Weight Decay 0.01
LR Schedule Cosine
The effect of # scales $K$.
9 More Experimental Results
Due to page limitations, the main paper reports only the strongest baseline methods in the primary comparison table. Here, we provide a more comprehensive evaluation that additionally includes 1-GAN [1-GAN] (GAN-based), PointFlow [PointFlow], DPF-Net [DPF-Net], SoftFlow [SoftFlow] (normalizing flow–based), SetVAE [setvae] (VAE-based), PVD-DDIM [DDIM] (diffusion models with advanced samplers), PSF [PSF] (flow-matching–based), and DIT-3D [DIT-3D]. We include only methods with publicly available implementations or those reporting explicit quantitative results in their original papers. Please see Table S3 for comprehensive comparisons.
Since ShapeNet categories are highly imbalanced, prior works typically report results only on the three largest categories—airplane, chair, and car—while the remaining categories contain significantly fewer samples. To provide a more comprehensive evaluation, we additionally report performance on three smaller categories: table, sofa, and lamp. As shown in Table S1, PointNSP consistently and substantially outperforms all baseline methods, demonstrating its strong learning efficiency under limited data conditions.
Model Table Sofa Lamp
CD / EMD CD / EMD CD / EMD
PointFlow [PointFlow] 92.3 / 92.8 85.6 / 88.4 88.7 / 98.5
PointGPT [PointGPT] 91.5 / 90.9 83.8 / 88.4 88.7 / 98.5
ShapeGF [ShapeGF] 83.7 / 81.5 79.2 / 81.3 86.4 / 86.8
PVD [PVD] 76.8 / 80.2 73.5 / 96.7 89.1 / 89.4
LION [LION] 68.4 / 78.7 67.9 / 90.2 81.3 / 82.6
PointNSP-m 61.2 / 76.5 60.8 / 83.7 73.9 / 74.8
Table S1: Per-category generation quality on ShapeNet. We report 1-NNA CD/EMD ($\downarrow$) on Table, Sofa, and Lamp. Best in bold, second best underlined.
In addition to the 55-class generation setting, we also evaluate the 13-class setting used in LION [LION]. The results, presented in Table S2, show that PointNSP achieves state-of-the-art generation quality.
Model CD $\downarrow$ EMD $\downarrow$
TreeGAN [TreeGAN] 96.80 96.60
PointFlow [PointFlow] 63.25 66.05
ShapeGF [ShapeGF] 55.65 59.00
SetVAE [setvae] 79.25 95.25
PDGN [PDGN] 71.05 86.00
DPF-Net [DPF-Net] 67.10 64.75
DPM [DPM] 62.30 86.50
PVD [PVD] 58.65 57.85
LION [LION] 55.52 53.89
PointNSP-m 54.70 52.82
Table S2: Generation results (1-NNA $\downarrow$) trained jointly on 13 classes of ShapeNet-vol.